OpenStudy (ray1998):

Algebra 2. Please help! Fan and medal will be given :) thank you! 1. Is y = 1/4 x^05 a power function? Explain your reasoning.

OpenStudy (anonymous):

Well I am pretty sure the answer is yes because power functions are in the form f(x)=kx^n.

OpenStudy (ray1998):

Okay, thank you :) can you give me another example so that I can better understand it?

OpenStudy (boldjon):

Quote{ a power function is a function of the form f(x) = cx^a, where c and a are constant real numbers and x is a variable. } in this case c = 1/4 , a=0.5 and they are constant real numbers

OpenStudy (ray1998):

That makes a lot of sense :) thank you! @boldjon

OpenStudy (boldjon):

ur welcome ^_^

OpenStudy (anonymous):

Another example could be f(x)=7x^2 7=k and 2=n (from my post above)